An electron is moving northward at 25 ,/s in a 4.5 T magnetic field pointing southward.

What is the force (magnitude and direction).
Answer #1

The force on a charged particle due to a magnetic field is given by the formula

F=q(TXB)

where q is the charge of the particle, \overrightarrow{v} is the velocity vector, and \overrightarrow{B} is the magnetic field vector.

Here, we know that the direction of the velocity is anti-parallel to the magnetic field, i.e they are oriented at 180^\circ to each other. You may already know that for vecotrs anti-parallel to each other, their cross product is zero. Here, we will use that result and prove the force on the electron is zero.

F=q(TXB)

From the definition of the cross product

F=g(TIB sinon)

where \overrightarrow{n} is a vector perpendicular to both the other vectors.

We know that Ꮎ = 180 . Substituting this, we get

F=g(TIB sin 180)

We also know that sin 180 = 0

Therefore, F =0 .

There is no force acting on the electron.